The combustion of ethane proceeds according to equation 2C2H6 + 7O2 → 4CO2 + 6H2O. The enthalpy change for combustion of 1 mol of this hydrocarbon equals –1600.0 kJ/mol. Use the following from Hess's law and calculate the enthalpy of formation of 1 mol of ethane, if ΔHf (CO2) = −393.5 kJ/mol, ΔHf (H2O) = −241.8 kJ/mol, ΔHf (O2) = 0.


1
Expert's answer
2022-05-04T09:21:03-0400

2C2H6 + 7O2 → 4CO2 + 6H2O, ΔH = –1600.0 kJ/mol


ΔHf(CO2) =−393.5 kJ/mol

ΔHf(O2) = 0 kJ/mol

ΔHf(H2O) = −241.8 kJ/mol


ΔHc = -2ΔH (C2H6) + 4ΔHf(CO2) + 6ΔHf(H2O) = -(–1600.0 kJ/mol) +4(−393.5 kJ/mol) + 6(−241.8 kJ/mol) = -1424.8 kJ/mol